# SQL-of-Thought: Multi-agentic Text-to-SQL with Guided Error Correction

### v1

@@ 1 Introduction
Text-to-SQL (NL2SQL) has emerged as a crucial problem in both research and real-world applications, enabling non-technical users to query structured databases through natural language. While early sequence-to-sequence and schema-aware models improved accessibility, they often struggled with generalization, resulting in subpar performance. Recent advances in large language models (LLMs) and prompting techniques have significantly raised performance, with methods such as DIN-SQL Pourreza and Rafiei (2023) and DAIL-SQL Gao et al. (2023) decomposing the task into subtasks. However, as highlighted by Li et al. Li et al. (2024) and Biswal et al. Biswal et al. (2024), even the most advanced systems remain brittle when dealing with realistic queries. This is because execution-based feedback alone cannot correct logically incorrect, yet syntactically valid, SQL queries.
Multi-agent approaches Wang et al. (2024); Pourreza et al. (2024); Shao et al. (2025) and reasoning-guided prompting Tai et al. (2023); Papicchio et al. (2025) have shown potential to bridge this gap. Multi-agent systems enhance modularity and specialization; however, their error correction still relies heavily on execution signals or static re-generation. On the other hand, reasoning-based methods improve query planning, yet unguided reasoning can often introduce new errors or inefficiencies. What is needed is a systematic way to combine multi-agent decomposition with structured reasoning and interpretable correction.
We tackle this issue using SQL-of-Thought, a multi-agent framework designed to enhance SQL queries. This framework incorporates several key components: (i) specialized agents for schema linking, identifying subproblems, generating query plans, synthesizing SQL, and correcting errors; (ii) a taxonomy-guided error correction loop that categorizes SQL failure modes and rectifies them through a Chain of Thought reasoning process; and (iii) an analysis comparing reasoning and non-reasoning models across different agent roles. Our framework achieves state-of-the-art execution accuracy on Spider Yu et al. (2019), Spider-Realistic Deng et al. (2021), and Spider-SYN Gan et al. (2021), while also showcasing effective cost-performance trade-offs through the use of hybrid models. These results demonstrate that compact, guided reasoning, paired with structured error feedback, is a more reliable approach for SQL querying than solely relying on execution-based refinement.
@@
  • State of the Art results on Spider Yu et al. (2019) benchmark and its variations: Spider-Realistic Deng et al. (2021), and Spider-SYN Gan et al. (2021).
@@ 2.1 History of NL2SQL
Initial works for Text-to-SQL used major sequence-to-sequence architectures such as PLMs (Programming Language Models), for example RESDSQL (Li et al., 2023), which identified relevant schema from the question and then used it to construct the SQL query. Other choices were GNNs (Graph Neural Networks), Recurrent Neural Networks (RNNs), Transformers, and most recently LLMs for natural language to SQL conversion. With the advent of LLMs, prompt engineering on powerful models like GPT (Brown et al., 2020; OpenAI et al., 2024) and Claude enabled fine-grained and faster solutions for NL2SQL. Shi et al (Shi et al., 2025) highlights how LLM-based solutions outperformed all other alternatives for programming text-to-SQL.
DIN-SQL (Pourreza and Rafiei, 2023) attempted in-context learning with error correction, decomposing subproblems, and defining different prompt templates to tackle each subtask. However, their method only regenerates prompts in case of failure, without specific error information. DAIL-SQL Gao et al. (2023) designs sophisticated SQL-code style prompt templates to encode the question and database schema. These templates provide few-shot examples for schema linking and SQL query generation, utilizing GPT-4 to generate solutions. DAILSQL-SC employs self-consistency to postprocess solutions for further refining.
Some studies Li et al. (2024) discuss production readiness and evaluation frameworks for NL2SQL using various prompting methods, again verifying that LLM-based approaches generally outperform pretrained task-specific models through higher reliability and generalization. They classify queries into four categories: subqueries, joins, logical connectors, and order-by clauses, to measure model efficacy across increasing levels of difficulty. The TAG framework Biswal et al. (2024) showed how existing models fail on complex real-world queries, highlighting that most NL2SQL models and benchmarks are only effective for roughly 20% of realistic user queries. They introduce a reasoning framework that decomposes the problem into three stages: (i) query synthesis through NL2SQL parsing, (ii) query execution over the underlying database, and (iii) answer generation, where an LLM integrates both the natural language query and execution results to produce a final human-readable answer.
The TAG framework Biswal et al. (2024) demonstrated how existing models struggle with complex real-world queries and proposed steps for query synthesis, execution, and answer generation. Multi-agent frameworks, such as Wang et al. Wang et al. (2025), introduced syntax validation and execution optimization, but relied mainly on execution feedback.
@@ 2.2 Agentic Approaches
Tool-SQL (Wang et al., 2024) pioneered the tool-first agentic framework, where an LLM-based agent iteratively generates and refines SQL queries with the aid of two tools: a database retriever to resolve condition mismatches and an error detector for stricter constraint validation. The approach primarily addressed database mismatches, such as missing highlights, altered paraphrasing, and similar issues.
Chase SQL (Pourreza et al., 2024) defines an elaborate framework consisting of value retrieval, multiple candidate generators and query fixers, and a selection agent. The candidates generate a query plan by prompting the LLM through an "EXPLAIN" keyword. The multi-agent framework proposed by Shao et al. (Shao et al., 2025) introduces a four-agent system where specialized roles (developer, researcher, executor, specialist) collaborate on simple and complex NL2SQL queries. By combining example retrieval, chain-of-thought reasoning, syntax validation, and error-driven refinement, the framework improves both correctness and execution efficiency.
@@ 2.3 Reasoning guided solutions to NL2SQL
ACT-SQL Zhang et al. (2023) experiments with prompting styles and few-shot exemplar selection strategies to automate prompt selection for natural language to SQL solutions, making the process cheap and time-saving. Think2SQL (Papicchio et al., 2025) and related methods explored reasoning for text-to-SQL by combining zero-shot prompting, supervised finetuning, and reinforcement learning. They asked whether reasoning improves text-to-SQL, how best to train reasoning-capable models, and whether execution-based rewards are sufficient. Their results show mixed evidence: reasoning models often perform better on query planning but not consistently across datasets. They also simplified schema linking by manually providing subsets, limiting the scope of evaluation. Tai et al. (Tai et al., 2023) highlight Chain-of-Thought prompting techniques, which achieve significant gains on the Spider dataset, and mention how unnecessarily detailed reasoning can propagate errors.
@@ 3.1 SQL-of-Thought
Query Plan Agent. The Query Plan Agent generates a step-by-step execution plan that maps the user’s intent to the schema and subproblems. Unlike prior work such as Chase-SQL Pourreza et al. (2024), which treats planning as a surface-level mapping, our agent is explicitly prompted to perform chain-of-thought reasoning to explain intermediate decisions. This structured reasoning encourages deeper alignment between the question, subproblems, and the final SQL query. The Query Plan Agent produces only the procedural plan and is explicitly restricted from generating executable SQL at this stage.
@@
Correction Plan Agent: The Correction Plan Agent initiates the correction loop by analyzing the failed SQL query in the context of the natural language question, schema, and execution results. Unlike systems such as DIN-SQL or DAIL-SQL Gao et al. (2023); Pourreza and Rafiei (2023), which rely solely on execution feedback, this agent is additionally guided by an error taxonomy derived from Shen et al. Shen et al. (2025). The taxonomy categorizes common error modes (e.g., schema mismatches, join inconsistencies, aggregation misuse), and the agent produces a chain-of-thought plan describing how to resolve the identified errors. Inspired by reflexive learning approaches Shinn et al. (2023), the agent iteratively generates structured feedback that informs the correction SQL generation process.
@@ 3.2 Error Taxonomy
The error taxonomy adopted in our correction loop, shown in Figure 2, is derived from the classification proposed in Shen et al. (2025), which systematically categorizes standard failure modes encountered in NL2SQL systems. Our proposed taxonomy expands on this by creating systematic categories of errors that can be identified and corrected by LLMs, as well as comprehensive subtypes for each. We provide concise error codes instead of lengthy explanations to facilitate easier identification and prevent overflowing the LLM context window. Our taxonomy spans a broad range of issues, including syntax errors (e.g., invalid aliases or malformed SQL), schema linking errors (e.g., missing or ambiguous columns, incorrect foreign keys), and join-related mistakes (such as missing joins, wrong join types, or inclusion of extra tables). It further covers errors in filter conditions (e.g., incorrect columns in the WHERE clause, type mismatches), aggregation logic (e.g., missing GROUP BY, misuse of HAVING), as well as value representation errors like hard-coded values or format mismatches. More advanced categories capture failures in subquery formulation (e.g., unused or incorrectly correlated subqueries), set operations (union, intersection, except), and other structural oversights, such as missing ORDER BY or LIMIT clauses. By explicitly codifying these error types, the taxonomy provides a fine-grained diagnostic lens that enables the correction plan agent to reason not just about what failed, but why it failed within the SQL generation pipeline, ranging from basic syntax issues to complex sub-query and join formulation errors, providing specific guidance for correction.
Incorporating this taxonomy into the correction loop enables the system to move beyond coarse execution-based feedback (as used in DIN-SQL Pourreza and Rafiei (2023) and DAIL-SQL Gao et al. (2023)) and instead provide interpretable, linguistically grounded guidance for iterative refinement. We prompt the correction plan agent with the summarized error taxonomy, along with a chain-of-thought reasoning template, which equips the agent to identify root causes of failure, align errors with schema or syntax constraints, and propose concrete repair strategies. This approach aligns with findings in reflexive learning approaches (Shinn et al., 2023), where LLMs benefit from verbalized self-feedback in enhancing decision-making for programming tasks. Thus, by leveraging the structured taxonomy from Shen et al. (2025), the multi-agent system establishes a principled loop of error detection, diagnosis, and guided correction, thereby closing the gap between raw execution feedback and systematic, explainable query refinement.
@@ 3.3.1 Dataset
We utilize the Spider dataset Yu et al. (2019) for evaluating our multi-agent framework. Spider offers a diverse range of database schemas and tasks, including 20 database settings and 1034 text-SQL pairs in its dev split. We avoid the BIRD-SQL benchmark due to the high quantity of annotation errors, as reported in Shen et al Shen et al. (2025). BIRD-SQL has fewer set operation-based questions than SPIDER, and also presents an evidence field for queries, which can add to hallucination while query generation. For these purposes, we test on Spider and its popular variant, Spider Realistic Deng et al. (2021), which is an evaluation set based on the Spider dev set with explicit mentions of column names removed, containing 508 samples. Spider SYN Gan et al. (2021) is a challenging variant of the Spider evaluation dataset. Spider- SYN is constructed by manually modifying natural language questions with synonym substitutions.
@@ 3.3.2 Evaluation Metrics
While we observe metrics such as Exact Match and Valid SQL Generation during our experiments, we only consider Execution Accuracy for holistic evaluation. One natural language question could have multiple SQL queries that answer it. LLMs frequently oversimplify, and in the case of Text-to-SQL tasks, they were found to assign variables to results from subqueries, which is why an Exact Match (EM) is not an accurate representation of the accuracy of the process. Thus, we adopt Execution Accuracy (EA) as the metric to evaluate our process. EA is a boolean field that compares the execution results of the generated query with the execution results of the Gold SQL label that exists in the dataset. Since the EA calculation requires the generated SQL query to be executed against Spider’s SQLite files, we also adopt SQLite SQLite for query generation.
@@ 3.4 Guided Error modification
Execution-based feedback alone is insufficient for error correction in NL2SQL. In our experiments, we observed that when using models such as Claude 3 Opus or GPT-4o-mini within the SQL-of-Thought framework, 95–99% of generated queries were already syntactically valid. Thus, traditional database execution errors were rare, and the predominant failures stemmed instead from intent mismatches, logically incorrect but syntactically valid queries, or errors involving incorrect or extraneous values through various join and limit clauses. Providing only execution traces Pourreza and Rafiei (2023); Gao et al. (2023) therefore offers little guidance for effective correction. To address this, we integrate the error taxonomy defined in Figure 2, which enumerates a wide range of SQL error types. By exposing the model to these structured error categories, the correction loop enables the agent to identify better and rectify subtle logical and semantic flaws. Empirically, on our strongest configuration (Claude 3 Opus + SQL-of-Thought), omitting the correction loop led to a 10% increase in incorrect queries compared to the whole system with guided error modification, as demonstrated through ablation on a 100-sample evaluation set in Table LABEL:different-model-results.
@@ 4 Results
We present the results of the proposed SQL-of-Thought framework on the Spider Yu et al. (2019) and Spider Realistic Deng et al. (2021) benchmarks (where results were available) in Table 1, highlighting the execution accuracy reported by earlier works and demonstrating how SQL-of-Thought outperforms existing methods. For all our experiments, we set the temperature to 0, and set top_p and other hyperparameters to default. These results are based on the Claude Opus 3 model, which serves as the foundation for evaluating our SQL-of-Thought framework. SQL-of-Thought reports an Execution Accuracy of 91.59% on Spider Yu et al. (2019) Benchmark and 90.16% Execution Accuracy on Spider Realistic Deng et al. (2021) benchmark. We also report an EA of 82.01% on Spider SYN Gan et al. (2021) benchmark. We also wish to report that for all our experiments, we observe a Valid SQL generation rate of 94%-99% with SQL-of-Thought.
@@
ChatGPT Brown et al. (2020) 74.4 -
GPT-4 (OpenAI 2023) OpenAI et al. (2024) 72.3 -
ACT-SQL + ChatGPT (Zhang et al. 2023) Zhang et al. (2023) 80.4 75.8
ACT-SQL + GPT-4 (Zhang et al. 2023) Zhang et al. (2023) 82.9 -
DIN-SQL + GPT-4 (Pourreza and Rafiei 2024) Pourreza and Rafiei (2023) 82.8 78.1
DAIL-SQL + GPT-4 (Gao et al. 2023) Gao et al. (2023) 83.1 75.6
DAIL-SQL + GPT-4 + SC (Gao et al. 2023) Gao et al. (2023) 83.6 75.2
MAC-SQL + GPT-4 (Wang et al. 2024) Wang et al. (2025) 86.8 -
Tool-SQL + GPT-4 (Wang et al., 2024) 86.9 82.9
Chase SQL (Pourreza et al., 2024) 87.6 -
@@
Table 2 denotes the results obtained when different models were employed as base LLMs in SQL-of-Thought on a 100-sample mini-batch from the Spider Dev dataset. We also evaluated the framework once with just the initial sequential flow, removing the self-correction loop, and once with direct SQL generation after Schema Linking and Subproblem Agents, removing the CoT Query Plan generation during both the initial and correction flows. The accuracy falls at least 5% without a Chain-of-Thought Query Plan before the SQL generation, and at least 8-10% for all models when the correction loop is not invoked. The models are listed in decreasing order of accuracy, noting that Claude Opus 3 performs the best, with GPT-5, GPT-4o-mini OpenAI et al. (2024), and finally GPT-3.5 Brown et al. (2020) producing 89%, 87% and 67% accuracy, respectively.

### v2

@@ 1 Introduction
Text-to-SQL (NL2SQL) has emerged as a crucial problem in both research and real-world applications, enabling non-technical users to query structured databases through natural language. While early sequence-to-sequence and schema-aware models improved accessibility, they often struggled with generalization, resulting in subpar performance. Recent advances in large language models (LLMs) and prompting techniques have significantly raised performance, with methods such as DIN-SQL [11] and DAIL-SQL [5] decomposing the task into subtasks. However, as highlighted by Li et al. [7] and Biswal et al. [1], even the most advanced systems remain brittle when dealing with realistic queries. This is because execution-based feedback alone cannot correct logically incorrect, yet syntactically valid, SQL queries.
Multi-agent approaches [21, 12, 14] and reasoning-guided prompting [19, 10] have shown potential to bridge this gap. Multi-agent systems enhance modularity and specialization; however, their error correction still relies heavily on execution signals or static re-generation. On the other hand, reasoning-based methods improve query planning, yet unguided reasoning can often introduce new errors or inefficiencies. What is needed is a systematic way to combine multi-agent decomposition with structured reasoning and interpretable correction.
We tackle this issue using SQL-of-Thought, a multi-agent framework designed to enhance SQL queries. This framework incorporates several key components: (i) specialized agents for schema linking, identifying subproblems, generating query plans, synthesizing SQL, and correcting errors; (ii) a taxonomy-guided error correction loop that categorizes SQL failure modes and rectifies them through a Chain of Thought reasoning process; and (iii) an analysis comparing reasoning and non-reasoning models across different agent roles. Our framework achieves state-of-the-art execution accuracy on Spider [22], Spider-Realistic [3], and Spider-SYN [4], while also showcasing effective cost-performance trade-offs through the use of hybrid models. These results demonstrate that compact, guided reasoning, paired with structured error feedback, is a more reliable approach for SQL querying than solely relying on execution-based refinement.
@@
  • State of the Art results on Spider [22] benchmark and its variations: Spider-Realistic [3], and Spider-SYN [4].
@@ 2.1 History of NL2SQL
Initial works for Text-to-SQL used major sequence-to-sequence architectures such as PLMs (Programming Language Models), for example RESDSQL [8], which identified relevant schema from the question and then used it to construct the SQL query. Other choices were GNNs (Graph Neural Networks), Recurrent Neural Networks (RNNs), Transformers, and most recently LLMs for natural language to SQL conversion. With the advent of LLMs, prompt engineering on powerful models like GPT [2, 9] and Claude enabled fine-grained and faster solutions for NL2SQL. Shi et al [16] highlights how LLM-based solutions outperformed all other alternatives for programming text-to-SQL.
DIN-SQL [11] attempted in-context learning with error correction, decomposing subproblems, and defining different prompt templates to tackle each subtask. However, their method only regenerates prompts in case of failure, without specific error information. DAIL-SQL [5] designs sophisticated SQL-code style prompt templates to encode the question and database schema. These templates provide few-shot examples for schema linking and SQL query generation, utilizing GPT-4 to generate solutions. DAILSQL-SC employs self-consistency to postprocess solutions for further refining.
Some studies [7] discuss production readiness and evaluation frameworks for NL2SQL using various prompting methods, again verifying that LLM-based approaches generally outperform pretrained task-specific models through higher reliability and generalization. They classify queries into four categories: subqueries, joins, logical connectors, and order-by clauses, to measure model efficacy across increasing levels of difficulty. The TAG framework [1] showed how existing models fail on complex real-world queries, highlighting that most NL2SQL models and benchmarks are only effective for roughly 20% of realistic user queries. They introduce a reasoning framework that decomposes the problem into three stages: (i) query synthesis through NL2SQL parsing, (ii) query execution over the underlying database, and (iii) answer generation, where an LLM integrates both the natural language query and execution results to produce a final human-readable answer.
The TAG framework [1] demonstrated how existing models struggle with complex real-world queries and proposed steps for query synthesis, execution, and answer generation. Multi-agent frameworks, such as Wang et al. [20], introduced syntax validation and execution optimization, but relied mainly on execution feedback.
@@ 2.2 Agentic Approaches
Tool-SQL [21] pioneered the tool-first agentic framework, where an LLM-based agent iteratively generates and refines SQL queries with the aid of two tools: a database retriever to resolve condition mismatches and an error detector for stricter constraint validation. The approach primarily addressed database mismatches, such as missing highlights, altered paraphrasing, and similar issues.
Chase SQL [12] defines an elaborate framework consisting of value retrieval, multiple candidate generators and query fixers, and a selection agent. The candidates generate a query plan by prompting the LLM through an "EXPLAIN" keyword. The multi-agent framework proposed by Shao et al. [14] introduces a four-agent system where specialized roles (developer, researcher, executor, specialist) collaborate on simple and complex NL2SQL queries. By combining example retrieval, chain-of-thought reasoning, syntax validation, and error-driven refinement, the framework improves both correctness and execution efficiency.
@@ 2.3 Reasoning guided solutions to NL2SQL
ACT-SQL [23] experiments with prompting styles and few-shot exemplar selection strategies to automate prompt selection for natural language to SQL solutions, making the process cheap and time-saving. Think2SQL [10] and related methods explored reasoning for text-to-SQL by combining zero-shot prompting, supervised finetuning, and reinforcement learning. They asked whether reasoning improves text-to-SQL, how best to train reasoning-capable models, and whether execution-based rewards are sufficient. Their results show mixed evidence: reasoning models often perform better on query planning but not consistently across datasets. They also simplified schema linking by manually providing subsets, limiting the scope of evaluation. Tai et al. [19] highlight Chain-of-Thought prompting techniques, which achieve significant gains on the Spider dataset, and mention how unnecessarily detailed reasoning can propagate errors.
@@ 3.1 SQL-of-Thought
Query Plan Agent. The Query Plan Agent generates a step-by-step execution plan that maps the user’s intent to the schema and subproblems. Unlike prior work such as Chase-SQL [12], which treats planning as a surface-level mapping, our agent is explicitly prompted to perform chain-of-thought reasoning to explain intermediate decisions. This structured reasoning encourages deeper alignment between the question, subproblems, and the final SQL query. The Query Plan Agent produces only the procedural plan and is explicitly restricted from generating executable SQL at this stage.
@@
Correction Plan Agent: The Correction Plan Agent initiates the correction loop by analyzing the failed SQL query in the context of the natural language question, schema, and execution results. Unlike systems such as DIN-SQL or DAIL-SQL [5, 11], which rely solely on execution feedback, this agent is additionally guided by an error taxonomy derived from Shen et al. [15]. The taxonomy categorizes common error modes (e.g., schema mismatches, join inconsistencies, aggregation misuse), and the agent produces a chain-of-thought plan describing how to resolve the identified errors. Inspired by reflexive learning approaches [17], the agent iteratively generates structured feedback that informs the correction SQL generation process.
@@ 3.2 Error Taxonomy
The error taxonomy adopted in our correction loop, shown in Figure 2, is derived from the classification proposed in Shen et al. [15], which systematically categorizes standard failure modes encountered in NL2SQL systems. Our proposed taxonomy expands on this by creating systematic categories of errors that can be identified and corrected by LLMs, as well as comprehensive subtypes for each. We provide concise error codes instead of lengthy explanations to facilitate easier identification and prevent overflowing the LLM context window. Our taxonomy spans a broad range of issues, including syntax errors (e.g., invalid aliases or malformed SQL), schema linking errors (e.g., missing or ambiguous columns, incorrect foreign keys), and join-related mistakes (such as missing joins, wrong join types, or inclusion of extra tables). It further covers errors in filter conditions (e.g., incorrect columns in the WHERE clause, type mismatches), aggregation logic (e.g., missing GROUP BY, misuse of HAVING), as well as value representation errors like hard-coded values or format mismatches. More advanced categories capture failures in subquery formulation (e.g., unused or incorrectly correlated subqueries), set operations (union, intersection, except), and other structural oversights, such as missing ORDER BY or LIMIT clauses. By explicitly codifying these error types, the taxonomy provides a fine-grained diagnostic lens that enables the correction plan agent to reason not just about what failed, but why it failed within the SQL generation pipeline, ranging from basic syntax issues to complex sub-query and join formulation errors, providing specific guidance for correction.
Incorporating this taxonomy into the correction loop enables the system to move beyond coarse execution-based feedback (as used in DIN-SQL [11] and DAIL-SQL [5]) and instead provide interpretable, linguistically grounded guidance for iterative refinement. We prompt the correction plan agent with the summarized error taxonomy, along with a chain-of-thought reasoning template, which equips the agent to identify root causes of failure, align errors with schema or syntax constraints, and propose concrete repair strategies. This approach aligns with findings in reflexive learning approaches [17], where LLMs benefit from verbalized self-feedback in enhancing decision-making for programming tasks. Thus, by leveraging the structured taxonomy from Shen et al. [15], the multi-agent system establishes a principled loop of error detection, diagnosis, and guided correction, thereby closing the gap between raw execution feedback and systematic, explainable query refinement.
@@ 3.3.1 Dataset
We utilize the Spider dataset [22] for evaluating our multi-agent framework. Spider offers a diverse range of database schemas and tasks, including 20 database settings and 1034 text-SQL pairs in its dev split. We avoid the BIRD-SQL benchmark due to the high quantity of annotation errors, as reported in Shen et al [15]. BIRD-SQL has fewer set operation-based questions than SPIDER, and also presents an evidence field for queries, which can add to hallucination while query generation. For these purposes, we test on Spider and its popular variant, Spider Realistic [3], which is an evaluation set based on the Spider dev set with explicit mentions of column names removed, containing 508 samples. Spider SYN [4] is a challenging variant of the Spider evaluation dataset. Spider- SYN is constructed by manually modifying natural language questions with synonym substitutions.
@@ 3.3.2 Evaluation Metrics
While we observe metrics such as Exact Match and Valid SQL Generation during our experiments, we only consider Execution Accuracy for holistic evaluation. One natural language question could have multiple SQL queries that answer it. LLMs frequently oversimplify, and in the case of Text-to-SQL tasks, they were found to assign variables to results from subqueries, which is why an Exact Match (EM) is not an accurate representation of the accuracy of the process. Thus, we adopt Execution Accuracy (EA) as the metric to evaluate our process. EA is a boolean field that compares the execution results of the generated query with the execution results of the Gold SQL label that exists in the dataset. Since the EA calculation requires the generated SQL query to be executed against Spider’s SQLite files, we also adopt SQLite [18] for query generation.
@@ 3.4 Guided Error modification
Execution-based feedback alone is insufficient for error correction in NL2SQL. In our experiments, we observed that when using models such as Claude 3 Opus or GPT-4o-mini within the SQL-of-Thought framework, 95–99% of generated queries were already syntactically valid. Thus, traditional database execution errors were rare, and the predominant failures stemmed instead from intent mismatches, logically incorrect but syntactically valid queries, or errors involving incorrect or extraneous values through various join and limit clauses. Providing only execution traces [11, 5] therefore offers little guidance for effective correction. To address this, we integrate the error taxonomy defined in Figure 2, which enumerates a wide range of SQL error types. By exposing the model to these structured error categories, the correction loop enables the agent to identify better and rectify subtle logical and semantic flaws. Empirically, on our strongest configuration (Claude 3 Opus + SQL-of-Thought), omitting the correction loop led to a 10% increase in incorrect queries compared to the whole system with guided error modification, as demonstrated through ablation on a 100-sample evaluation set in Table LABEL:different-model-results.
@@ 4 Results
We present the results of the proposed SQL-of-Thought framework on the Spider [22] and Spider Realistic [3] benchmarks (where results were available) in Table 1, highlighting the execution accuracy reported by earlier works and demonstrating how SQL-of-Thought outperforms existing methods. For all our experiments, we set the temperature to 0, and set top_p and other hyperparameters to default. These results are based on the Claude Opus 3 model, which serves as the foundation for evaluating our SQL-of-Thought framework. SQL-of-Thought reports an Execution Accuracy of 91.59% on Spider [22] Benchmark and 90.16% Execution Accuracy on Spider Realistic [3] benchmark. We also report an EA of 82.01% on Spider SYN [4] benchmark. We also wish to report that for all our experiments, we observe a Valid SQL generation rate of 94%-99% with SQL-of-Thought.
@@
ChatGPT [2] 74.4 -
GPT-4 (OpenAI 2023) [9] 72.3 -
ACT-SQL + ChatGPT (Zhang et al. 2023) [23] 80.4 75.8
ACT-SQL + GPT-4 (Zhang et al. 2023) [23] 82.9 -
DIN-SQL + GPT-4 (Pourreza and Rafiei 2024) [11] 82.8 78.1
DAIL-SQL + GPT-4 (Gao et al. 2023) [5] 83.1 75.6
DAIL-SQL + GPT-4 + SC (Gao et al. 2023) [5] 83.6 75.2
MAC-SQL + GPT-4 (Wang et al. 2024) [20] 86.8 -
Tool-SQL + GPT-4 [21] 86.9 82.9
Chase SQL [12] 87.6 -
@@
Table 2 denotes the results obtained when different models were employed as base LLMs in SQL-of-Thought on a 100-sample mini-batch from the Spider Dev dataset. We also evaluated the framework once with just the initial sequential flow, removing the self-correction loop, and once with direct SQL generation after Schema Linking and Subproblem Agents, removing the CoT Query Plan generation during both the initial and correction flows. The accuracy falls at least 5% without a Chain-of-Thought Query Plan before the SQL generation, and at least 8-10% for all models when the correction loop is not invoked. The models are listed in decreasing order of accuracy, noting that Claude Opus 3 performs the best, with GPT-5, GPT-4o-mini [9], and finally GPT-3.5 [2] producing 89%, 87% and 67% accuracy, respectively.